OpenStudy (anonymous):

A man is 20 older than his son. In ten years, he will be twice as old as his son will then be. How old is each?

OpenStudy (anonymous):

please show your solution.

OpenStudy (anonymous):

in ten years : 20 + 10 = 30 /2 = 15 now : 30-10= 20 15-10=5 boy is 5 father is 20

OpenStudy (anonymous):

the / is a division symbol

OpenStudy (anonymous):